OpenStudy (anonymous):

12 + 0.35x = 20.05 A. 91.5 B. 57.3 C. 2.8175 D. 23

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (doggy):

Simplifying 12 + 0.35x = 20.05 Solving 12 + 0.35x = 20.05 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-12' to each side of the equation. 12 + -12 + 0.35x = 20.05 + -12 Combine like terms: 12 + -12 = 0 0 + 0.35x = 20.05 + -12 0.35x = 20.05 + -12 Combine like terms: 20.05 + -12 = 8.05 0.35x = 8.05 Divide each side by '0.35'. x = 23 Simplifying x = 23

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thank you guys! :)